Based on the corporate valuation model, the value of a company’s operations is $1,200 million. The company’s balance sheet shows $80 million in accounts receivable, $60 million in inventory, and $100 million in short-term investments that are unrelated to operations. The balance sheet also shows $90 million in accounts payable, $120 million in notes payable, $300 million in long-term debt, $50 million in preferred stock, $180 million in retained earnings, and $800 million in total common equity. If the company has 30 million shares of stock outstanding, what is the best estimate of the stock’s price per share?
a. $24.90
b. $27.67
c. $30.43
d. $33.48
e. $36.82
1
Expert's answer
2012-08-17T08:17:35-0400
b. $27.67
Market Value of company = 1200M + 100M = 1300M
MV of company = MV of debt + MV of preferred + MV of equity = 420M + 50M + MV Equity
=> MV Equity = 1300M - 420M - 50M = 830M (assume book value of debt approximately = MV debt)
=> share price = 830/30 = $27.67